The patel manufacturing company allows for a defect rate of 0.9% on its production line.The defect average for last week was 3 times out of 1,500.What is the difference between the allowed defect rate and the average for last week?: The patel manufacturing company allows for a defect rate of 0.9% on its production line.The defect average for last week was 3 times out of 1,500.What is the difference between the allowed defect rate and the average for last week?
Answer by Fombitz(1756)

3/1500=0.002
In terms of percentage that would be,
0.002=0.2/100 or 0.2%.
The difference is
D=0.9-0.2=0.7 or 0.7% better than average.